OpenStudy (anonymous):

A pure resistance and a pure inductance are connected in series across a 100V (rms) AC line. An AC voltmeter gives the same reading when connected across either the R or the L. What does it read? (note: meters read rms values)

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

Sorry I'm kinda busy right now

OpenStudy (anonymous):

it's okay

OpenStudy (kropot72):

The equal voltages are 90 degrees out of phase.  \[\cos 45\deg=\frac{V _{R}}{100}\] Can you find the voltmeter reading now?

OpenStudy (kropot72):

@Lynncake Are you there?

OpenStudy (anonymous):

@kropot72 i got the value to be 70.71 Is this right? and why is the angle 45 degrees?

OpenStudy (kropot72):

Yes the voltages across the R and the L are equal with value \[\frac{100}{\sqrt{2}}\] In this case the applied voltage is the vector sum of the voltages across the R and the L and has the phase angle 45 degrees with respect to the circuit current.

OpenStudy (anonymous):

So, if the magnitudes of R and L are altered so that the voltmeter across L reads 50V. The voltmeter that is read across R is 86.6 \[\sqrt{100^2 - (50)^}\]